If the average of 27-x, x-8, and 3x+11is y, what is the average of 2y and 2y over 3?
Answer by Boreal(15235)

27-x +x-8+3x+11=3x+30=3(x+10)
average is (3/3)(x+10) =
x+10
y=x+10
2y= 2x+20
(2y/3) = (2x+20)/3
The average of those is (1/2)[ (2x+20) +(2x+20/3]
Common denominator of 3 for what is inside parentheses
(1/2)(1/3)[6x+60+2x+20]
(1/6) (8x + 80)
(4/3) (x + 10)
average of 2y and 2y/3 = 8y/3/2= (4/3) y
but y= x+10
(4/3)y= (4/3) (x+10)